OpenStudy (anonymous):

What is the vertex of the graph of y = –4(x + 2)2 + 5? (2, 5) (–2, 5) (5, –2) (5, 2) Solve. 5x2 – 20 = 0 x = 5 and x = 4 x = 4 and x = 2 x = –5 and x = 20 x = 2 and x = –2 I think the first one is A and the second one is C

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Callisto

OpenStudy (aripotta):

the vertex is (-h, k) isn't it? so the vertex should be (-2, 5)

OpenStudy (phi):

look at this http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php can you figure it out ?

OpenStudy (anonymous):

ah no

OpenStudy (aripotta):

5x^2 – 20 = 0 5x^2 = 20 x^2 = 4 sqrt(x^2) = +/- sqrt(4) x = +/- 2

OpenStudy (aripotta):

so the second one is d

OpenStudy (aripotta):

does that make sense?

OpenStudy (anonymous):

yes.

OpenStudy (aripotta):

ok :]

OpenStudy (anonymous):

what about the first one ?

OpenStudy (aripotta):

look at my first post